Status Report

ISS On-Orbit Status 31 Jan 2003

By SpaceRef Editor
January 31, 2003
Filed under , ,
ISS On-Orbit Status 31 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.

Wake-up was at the regular 1:00am EST, followed by post-sleep activities and breakfast.  Begin of sleep time: 4:30pm.

FE-1 Nikolai Budarin and CDR Kenneth Bowersox continued working on the Progress 9P spacecraft in preparation for its undocking tomorrow morning.  After power-down of the cargo vehicle’s US-21 matching unit by MCC-M/TsUP and disconnection of the BITS 2-12 onboard telemetry measurement system from it by Budarin, he and Bowersox removed the no-longer-needed US-21 in its container box.  [The matching unit and associated commutator gear provide the electronic interface between the Service Module (SM) and Progress for SM computer control of Progress propulsion, i.e., during ISS reboost maneuvers.  The avionics will be recycled on a future cargo flight.  Progress undocking and post-separation control will be handled by the drone’s pre-programmed automated sequencing system.]

After the R&R, Budarin and Bowersox conducted final undocking preparations.  Separation command will nominally be given tomorrow morning at 10:59am EST, followed by deorbit burn ignition at 2:10pm and atmospheric entry beginning at 2:47pm.  Plume observations will be conducted from the SM with the infrared spectrometer of the Russian geophysical GFI-1 “Relaksatsiya” experiment.  [Nikolai activated Progress and removed the air duct between it and the SM transfer tunnel (PrK) at Zvezda’s aft end.  QD (quick disconnect) screw clamps were removed and the Progress/SM interface was video-surveyed with the U.S. Sony PD-100 DVCAM for subsequent downlink to and inspection by TsUP.  Hatch closure proceeded at about 9:30am EST (which also served as decision point for final irreversible launch preparations at Baikonur for Progress 10P), followed by an interface leak check of the Progress/PrK vestibule, during which Russian thrusters were disabled.]

Launch countdown for Progress 10P (M-247) at Baikonur continues toward liftoff on 2/2 (Sunday), at 7:59am EST.  Docking is set for 2/4 (Tuesday), at 9:48am.  [Since the ISS solar arrays will be feathered during the 2-hr. docking window (i.e., not tracking the Sun), the expected power generation decrease will require some temporary powerdowns in the U.S. segment (USOS) and of one USOS-to-RS (Russian segment) SNT-50MP voltage and current stabilizer unit (transformer).  Details are still in work.]

As POC (Payload Operations Center) continues to troubleshoot the recent failure of the ER3 ARIS (EXPRESS rack 3/active rack isolation system) calibration, Bowersox continued ER work, today loading the ER2 laptop (ELC) with the new software, but not yet the RIC (rack interface controller).  [The new load was expected to help troubleshoot why the laptop is not communicating with the rack.  If it is due to a corrupt file, as suspected, the s/w load will fix the problem.  If it is a hardware problem, however, the load may not complete successfully.  Prior to the load, Don Pettit needed to swap out the ELC’s hard drive with a spare (with Windows NT, instead of Windows 95 or 98).]

Today, the highly successful EarthKAM operation came to an end for Increment 6.  FE-2/SO Pettit deactivated the camera system and stowed it, reconfiguring the cabling to get the Lab window ready for the installation of the WORF (window observational research facility), a modified ISPR (international standard payload rack) due to arrive in early March on STS-114/ULF-1.

Budarin checked up on the MO-21 “Ecosfera” air sampler and incubation equipment, monitoring colony growth for atmospheric microbial analysis.  [MO-21 determines microbial contamination of the ISS atmosphere, specifically the total bacterial and fungal microflora counts and microflora composition according to morphologic criteria of microorganism colonies.]

Later, he was scheduled to tag up with an Uragan (“hurricane”) specialist to discuss the new earth imaging sessions planned for tomorrow and Sunday.

After completion of the regeneration process on the second Metox (metal oxide) canister, which can now be used again in the next EVA operations as a carbon dioxide (CO2) filter, the crew monitored the deactivation of the U.S. CDRA (CO2 removal assembly).   The ground also reset the Airlock temperature setpoint back up to 70F (21 degC).

In the RS, the Vozdukh CO2 scrubber remains powered off for a long-planned efficiency test of the Russian LiOH (lithium hydroxide) cartridges.  [Intended to support extension of certified life time of the backup CO2 absorbers, the activation of one cartridge was to test its capability to remove CO2 from the cabin without Vozdukh or CDRA running.]

Recent ppCO2 (CO2 partial pressure) values have consistently been higher in the USOS than in the SM, as indicated by the Russian sensor system.  For the LiOH canister efficiency test in the RS, it was suggested to Moscow to use the portable U.S. CDMK (CO2 monitoring kit) instrument instead, to ensure that Lab ppCO2 concentration does not go over the level of 6.0 mmHg.

CDR Bowersox set up the video equipment for recording the crew’s workout on the RED (resistive exercise device).  All three crewmembers were scheduled to record their anaerobic workout, and the video camera was to be stowed after the last session of the day.  Later, Don Pettit performed the regular weekly downloading of exercise files from the TVIS treadmill, RED and CEVIS bike to the MEC (medical equipment computer).

Nikolai Budarin completed his daily checkup of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) plant growth experiment.

Don Pettit performed the regular daily tasks of SOSh life support systems maintenance and IMS database delta file preparation, while Ken Bowersox did the Lab payload status checkup (PCG-STES010).

Sox also began the first part of an extensive “pantry audit” of on-board provisions, entailing such item categories as office supplies, towels and napkins, printer supplies, batteries, sanitary hygiene provisions, and hygiene skip-cycle supply (in case of delayed resupply).  Due to Increment 7/ULF-1 packing deadlines, today’s audit gave priority to the office supplies part.  [The audit is intended to verify the complement of U.S. crew provisions onboard, to allow MCC-H to determine necessary resupply items and which items are not being used and are not needed, and to support Increment 7/ULF-1 packing as well as Increment 8 manifesting.]

Pending final signature of the jointly prepared water venting protocol by the Russian side, the next Lab condensate dump will take place next Friday (2/7), performed non-propulsively (i.e., from two vents) in the current LVLH attitude, as opposed to the more regularly used XPOP attitude.  If the agreement is not signed in time, the excess condensate will be dumped into a CWC (contingency water container).

Troubleshooting of the tripped RPCM (remote power controller module) continues.  If the VTR1 (video tape recorder #1), which it powers, is found not at fault, the RPCM will be changed out.  If the VTR is the cause of the trip, there are two VTR spares on the ground, ready to go if needed to replace it (typically on 12A, but this may change).

As reported before, the four valves and the electronics inside the P1 ETCS NTA (external thermal control system/nitrogen tank assembly) with boith Loop B heaters failed are operating well.  Analysis is now concentrating on determining the temperatures to be expected for the NTA over the next few weeks.  Preliminary data indicate that keeping the interior electronics operating may generate enough heat to prevent hardware problems.

Readiness review for the new C&C (command & control) software upgrade R.3 will be conducted at JSC on 2/5 (Wednesday), with Russian participation.  Prior to the R.3 upload, Moscow will upload the new Russian patch 5.04, followed by the 7.01 patch after the R.3 upload. 

STS-107/Columbia is scheduled to land at KSC tomorrow (Saturday) at 9:16am EST.

Today’s targets for the CEO (crew earth observations program) were Irrawaddy River Delta, Burma (nadir pass.  Sun glint opportunity left of track on inner delta waterways), Rangoon, Burma (nadir pass.  The city lies on an east-side arm of the delta), Bombay, India (a touch left of track.  The city lies on the one major bay on India’s otherwise linear west coast.  Oblique views left may reveal aerosol masses related to this large city), Nairobi, Kenya (nadir pass; ESC [electronic still camera]), Patagonian Glaciers (southern end of the twin ice fields is likely to remain clear of cloud.  Crew was to try for glacier tongues on both sides of the Andes, especially the Pacific side where coverage is poorer), Central Africa (Dynamic event. Equatorial cloud masses lie further south over Africa at the height of “southern summer” than in any other part of the globe.  Numerous fires are raging in the Sahel savannas.  Looking obliquely left and right of track.  Also, there was an opportunity to document enigmatic river patterns in the vast Salamat swampland of the Central African Republic), West Africa Fires (Dynamic event. Numerous fires are raging in the Sahel.  Looking obliquely left and right of track.  Oblique images of the coastline and rainforest-savanna boundary were also requested under present clear skies), Panama panorama (Dynamic event.  Clear winter skies should allow a left-looking view north up the central American isthmus.  Requested was a long panorama in this seldom imaged part of the world), and Puerto Rico (fringing coral reef mapping images were requested for near nadir reefs on the west end of the island).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32 Amp mode).  Vozdukh CO2 scrubber is Off.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 753; temperature (deg C) — 23.1; ppO2 (mmHg) — 160.4; ppCO2 (mmHg) — 2.5 (suspect).
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 19.1.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 738.16; temperature (deg C) — 22.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 740.27; temperature (deg C) — 22.7; ppO2 (mmHg) — 162.9; ppCO2 (mmHg) — 3.4;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 740.37; temperature (deg C) — 20.0; shell heater temp (deg C) — 20.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.7
  • PMA-2:  Shell heater temp (deg C) — 18.4

(n/a = data not available)

Propulsion System (PS): 

  • Total propellant load available [SM(820) + FGB(2898) + Progress(0)] — 3718 kg (8197 lb) as of 1/30/03.  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B both in “blind” dual-angle mode (directed position).
  • SM batteries:  Battery #7 is off line (failed); battery #2  is offline (ROM/capacity restoration mode); all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries:  Battery #2 is offline; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-1 is On (primary), EXT-2 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is out of the set (as of 11/14/02).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1 (GPS)
  • Attitude — U.S. SIGI-1 (GPS)
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • LVLH TEA (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.1 deg, roll: 0 deg]), CMG/TA (thruster assist) momentum management).
  • Solar Beta Angle:  12.2 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally, but VTR1 is off.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched at WS4, with KA power. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:40am EST [= epoch]):

  • Mean altitude — 387.6 km
  • Apogee — 391.4 km
  • Perigee — 383.9 km
  • Period — 92.31 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0005582
  • Orbits per 24-hr. day — 15.60
  • Mean altitude loss in last 24 hours — 200 m
  • Solar Beta Angle — 12.2 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 23968
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.